# Observation of Cosmic-Ray Anisotropy in the Southern Hemisphere with Twelve Years of Data Collected by the IceCube Neutrino Observatory

### v1

@@ Abstract
We analyzed the 7.92×$\times 10^{11}$ cosmic-ray-induced muon events collected by the IceCube Neutrino Observatory from May 13, 2011, when the fully constructed experiment started to take data, to May 12, 2023. This dataset provides an up-to-date cosmic-ray arrival direction distribution in the Southern Hemisphere with unprecedented statistical accuracy covering more than a full period length of a solar cycle. Improvements in Monte Carlo event simulation and better handling of year-to-year differences in data processing significantly reduce systematic uncertainties below the level of statistical fluctuations compared to the previously published results. We confirm the observation of a change in the angular structure of the cosmic-ray anisotropy between 10 TeV and 1 PeV, more specifically in the 100–300 TeV energy range. For the first time, we analyzed the angular power spectrum at different energies. The observed variations of the power spectra with energy suggest relatively reduced large-scale features at high energy compared to those of medium and small scales. The large volume of data enhances the statistical significance at higher energies, up to the PeV scale, and smaller angular scales, down to approximately 6 compared to previous findings.
@@ 1 Introduction
It is widely believed that most cosmic-ray particles detected on Earth below EeV energy are produced within the Milky Way galaxy. Their energy spectrum and arrival directions are determined by the properties and distribution of their galactic sources and by their propagation through the turbulent and inhomogeneous interstellar magnetized plasma. Although cosmic-ray properties have been investigated for over a century, it was only in the last thirty years that extensive ground-based experiments have accumulated high-quality observations of the arrival direction distribution of cosmic rays (Nagashima et al., 1998; Hall et al., 1999; Guillian et al., 2007; Aglietta et al., 2009; Abdo et al., 2009; Munakata et al., 2010; De Jong, 2011; Aartsen et al., 2016a; Amenomori et al., 2017; Bartoli et al., 2018; Abeysekara et al., 2018, 2019; Apel et al., 2019; Ahlers, 2019; He, 2023).
@@
In this paper, we report the cosmic-ray anisotropy results obtained by the IceCube Neutrino Observatory, with 7.92×$\times 10^{11}$ cosmic-ray events collected over a period of twelve years. In Sec. 2, we describe the IceCube experiment, including the data used in this analysis and the simulation used to estimate the cosmic-ray energy of each event in the sample. In Sec. 3, we describe how the analysis is performed, what methods and tools we used, and the results. We then describe how we estimated the systematic uncertainties associated with the anisotropy measurement in Sec. 4 and conclude with a summary discussion in Sec. 5 along with an outlook in Sec. 6 on how this observation can contribute to our understanding of cosmic-ray physics and what additional measurements are necessary.
@@ 2.1 Detector
The IceCube Neutrino Observatory (Aartsen et al., 2017) is a cubic-kilometer-scale detector located at the geographic South Pole. The largest component of the detector is an in-ice array buried deep in the Antarctic ice shelf, consisting of 5,160 digital optical modules (DOMs), extending from a depth of 1,450 meters to about 2,450 meters. The in-ice array is augmented with a denser subarray called DeepCore (Abbasi et al., 2012a). IceCube also includes a surface array called IceTop (Abbasi et al., 2013), composed of tanks with DOMs inserted in transparent ice and designed to detect cosmic-ray showers above 100 TeV. This paper describes an investigation performed with the high-energy cosmic-ray muon data collected by the IceCube in-ice array.
@@
The IceCube in-ice array is designed to detect high-energy neutrinos of astrophysical origin. Neutrinos are observed through the secondary charged particles generated by their interaction in the ice inside and outside the instrumented volume or in the bedrock below. Such secondary particles, being relativistic, induce Cherenkov photons that propagate through the ice and are detected by the DOMs. The PMT produces a current pulse whenever a Cherenkov photon ejects an electron from its cathode; these pulses are digitized and time-stamped by the DOM onboard electronics before being sent to a data acquisition system for trigger determination and subsequent event reconstruction. To identify a physical event and filter out background signals caused by the 500-Hz PMTs’ dark noise, neighboring DOMs on the same string must satisfy a local coincidence in time within ±$\pm 1\,\mu$s. A trigger is issued when eight or more DOMs satisfying the local coincidence criterion record photons within a ±$\pm 5\,\mu$s time window (Abbasi et al., 2009; Aartsen et al., 2017).
All events that trigger the IceCube detector are reconstructed using a likelihood-based method that accounts for light propagation in the ice. This Single-Photo-Electron (SPE) reconstruction maximizes the likelihood constructed from the probability density function for the arrival times of single photons at the locations of activated optical modules to reconstruct the muon track (Ahrens et al., 2004). This SPE likelihood algorithm is seeded with the results of a fast linefit reconstruction that minimizes the sum of the squares of the distances between the track and the recorded pulses at each optical module (Aartsen et al., 2014).
@@ 2.2 Experimental Data
This analysis uses twelve full years of data collected by the completed IceCube detector (referred to as IC86) from May 13, 2011, through May 12, 2023. The data correspond to the good run list established by IceCube’s online monitoring system (Aartsen et al., 2017).
The rate of events collected by the IceCube in-ice array used in this work is between 2.0 and 2.3 kHz, following the yearly modulation of the stratospheric temperature over the Antarctic continent (Tilav et al., 2019; Gaisser & Verpoest, 2021; Verpoest et al., 2024). The volume of data collected amounts to about 1 TB per day. The limited bandwidth of the communication satellites that provide connectivity to the South Pole sets a practical limit on the amount of information that can be stored and transferred from each event that activates the detector’s trigger. For this reason, we utilize a compact data storage and transfer (DST) format to store the reconstruction results of all recorded events performed in situ before transferring the data to the northern hemisphere.
@@ 2.3 Simulation Data
We propagate the muons generated by CORSIKA through the ice and simulate the in-ice array response. The detector simulation — including detailed photon propagation, PMT response, and acquisition electronics — uses IceCube’s IceTray framework (DeYoung, 2004) with an interface to read CORSIKA formatted files. This simulation uses only high-energy muons reaching the in-ice array; all other shower components not contributing to the physical event are omitted to reduce memory requirements and computation time. Finally, the triggered muon events are reconstructed and processed with the same pipeline used for experimental data.
@@ 3.1 Constructing Arrival Direction Sky Maps
Numerous methods are used to measure the anisotropies in cosmic-ray arrival direction, depending on the experiment being used, including: Rayleigh amplitude method (Linsley, 1975; Antoni et al., 2004), time-scrambling (Alexandreas et al., 1993), direct integration (Atkins et al., 2003), a $\chi^{2}$ iterative method optimized for gamma-ray searches (Amenomori et al., 2005), East-West (Bonino et al., 2011), and the iterative maximum likelihood method developed by Ahlers et al. (2016). This last method was utilized in the combined analysis with HAWC in Abeysekara et al. (2019). It is designed to correctly recover the amplitude of large-scale features in the anisotropy that are attenuated as a result of the difference between the instantaneous and integrated field of view (FoV). For detectors located away from the poles, the instantaneous exposure varies during a 24 hr period and does not match the full daily exposure. However, as noted by Ahlers et al. (2016), since IceCube’s instantaneous FoV is identical to the time-integrated one, the iterative method is equivalent to direct integration or time scrambling. For this reason, in this analysis, we employed the time-scrambling method, as in previous IceCube analyses (Abbasi et al., 2012b; Aartsen et al., 2016a), to provide a “reference sky map” representing the experiment’s response to an isotropic cosmic-ray flux. This reference sky map also accounts for possible detection biases (e.g., nonuniform exposure and gaps in the experiment’s livetime).
@@
where $N_{i}$ and $\langle N\rangle_{i}$ represent the number of events for pixel $i$ in the data and reference sky maps, respectively. The statistical significance of each pixel is estimated using the method developed by Li & Ma (1983). We smooth the data and the reference sky maps for visualization purposes to observe large- and small-scale anisotropies at high significance (see Sec. 3.3). We use a process known as top-hat smoothing, which assigns a value in each pixel equal to the sum of all pixels’ content within a smoothing radius. Our default smoothing radius is $20^{\circ}$, consistent with previous IceCube publications (Abbasi et al., 2011, 2012b; Aartsen et al., 2016a). Analyses of the cosmic-ray arrival direction distributions are conducted on unsmoothed data and reference maps.
The Monte Carlo simulation data described in Sec. 2.3 indicate that the median angular resolution of the entire data sample is about $3^{\circ}$ (reaching $1^{\circ}$ above 100 TeV). However, it surpasses $20^{\circ}$ above a zenith angle of $50^{\circ}$ at 10 TeV, with a significant improvement at higher energy. The poor angular resolution at large zenith angles results from the increasing energy threshold of cosmic-ray primaries, as muons must travel increasingly longer distances through the ice. As a result, the number of hit optical sensors decreases at larger zenith angles for a given primary energy (Abbasi et al., 2012b; Abeysekara et al., 2019). For visualization purposes, sky maps are masked to omit events with reconstructed zenith angles above $65^{\circ}$, above which the event rate is negligible. On the other hand, smaller sky coverage is used for quantitative data analyses (see Sec. 3.4). This study does not attempt to select events with better angular resolution, as we aim to obtain the largest data volume to determine anisotropy structures down to approximately $6^{\circ}$ (see Sec. 3.5).
@@ 3.2 Constructing Energy Datasets
To study the energy dependence of the cosmic-ray anisotropy, we split the experimental data into several sets of increasing cosmic-ray energy. This procedure uses the simulation data described in Sec. 2.3. In earlier studies on anisotropy, we relied on the data gathered during the construction of IceCube, which limited our capacity to generate sufficient simulation data for each experiment’s configuration (see Aartsen et al. (2016a) and references therein). However, we now benefit from exclusively using the final detector configuration for the entire time period used in this work.
@@
The primary cosmic ray energy estimation is done on a statistical basis using the number of DOMs triggered by each event ($N_{\mathrm{channel}}$) and the corresponding muon trajectory’s reconstructed zenith angle ($\theta_{\mathrm{reco}}$). Higher deposited energy in the experiment triggers more DOMs. This happens with high-energy muons and high-multiplicity muon bundles, which are associated with more energetic or heavier cosmic-ray primary particles. In addition, events with a larger zenith angle must penetrate more ice before passing through IceCube’s in-ice array. Inclined muons therefore have higher mean energy at the surface, corresponding to higher primary cosmic-ray energy (Aartsen et al., 2016b). We use the simulation data described in Sec. 2.3 to determine the dependency between the experiment’s response and the cosmic-ray primary particle energy as a function of the reconstructed zenith angle.
@@
Using the procedure described in Aartsen et al. (2016a), Monte Carlo simulated events were placed into a three-dimensional histogram with axes corresponding to the observed $\log_{10}(N_{\mathrm{channel}})$, $\cos(\theta_{\mathrm{reco}})$, and the logarithm of the true energy of the primary cosmic-ray particle (bin sizes of 0.044, 0.07, and 0.05, respectively). After the histogram is populated, we select the median true energy value for each bin, effectively transforming the histogram into a lookup table that displays median energy as a function of $\log_{10}(N_{\mathrm{channel}})$ and $\cos(\theta_{\mathrm{reco}})$. The table is smoothed with a B-spline function (Whitehorn et al., 2013) to avoid statistical artifacts resulting from the limited simulation at high energies. The end result is used to define nine ranges of median primary cosmic-ray energy, as shown in Fig. 1a. The bands in the figure provide a proxy for estimating the primary energy of events, based on the measured variables $N_{\mathrm{channel}}$ and $\theta_{\mathrm{reco}}$. The described energy selection method yields an energy distribution in each event set defined by a colored band with a median value that is fairly constant as a function of $\theta_{reco}$. To ensure that the median energy on the lowest energy bin is also constant as a function of zenith angle, we exclude the events populating the white area in the bottom corner of Fig. 1a. Since the anisotropy changes with energy (see Sec. 3.3), ensuring a uniform energy response across the FoV provides an unbiased measurement of the arrival direction distribution.
@@
Table 2 displays the number of events selected for each energy set used in the analysis. This corresponds to about 82% of the total number of events collected by the IceCube in-ice array (see Table 1), a reduction stemming from the limited FoV (see Sec. 3.1), the exclusion of the lower right corner ($\log_{10}(N_{\mathrm{channel}})$, $\cos(\theta_{\mathrm{reco}})$) region in Fig. 1a, and the selection $N_{\mathrm{channel}}>10$, which removes events that are harder to reconstruct. For all energy bins, we apply a reconstructed zenith angle cut of $65\degree$. The only exception is the lowest energy bin, at $E_{\mathrm{median}}=13$ TeV, where we apply a reconstructed zenith angle cut of $60\degree$, since for larger zenith angles, events fall below the $N_{channel}>10$ threshold. For the highest energy bin, we apply only a lower bound in the energy selection. The primary cosmic-ray energy distributions corresponding to each of the nine selected regions are shown in Fig. 1b, along with the associated median energy. Note that the energy distributions of these sets overlap significantly. However, the energy sets are statistically independent as each event is placed in only one of them. The overlap between the different distributions results mainly from two reasons. First, as in the case of the angular resolution (see Sec. 3.1), there is no attempt to optimize the muon energy resolution with dedicated event selections. The goal is to have the largest number of events possible while retaining a reasonable quality. Secondly, the energy deposited by the muon bundles has wide event-by-event fluctuations, and even with a more accurate muon energy estimation, its relation to the parent cosmic-ray particles would still be affected by large fluctuations.
By comparing the H3a (Gaisser, 2012), Polygonato (Jörg R. Hörandel, 2003), and GST (Gaisser et al., 2013) composition models, we determined that the uncertainty in the median energy from different cosmic-ray flux models is less than $4$ in $\log_{10}(E/\mathrm{GeV})$. The predicted atmospheric muon fluxes depend on the assumed primary cosmic-ray spectrum and composition and the hadronic models that describe these interactions. Discrepancies among these models lead to an overall value of 20–40% in the predicted muon flux within the energy range from 1 TeV to 1 PeV (Riehn et al., 2020).

### v2

# Observation of Cosmic-Ray Anisotropy in the Southern Hemisphere with 12 yr of Data Collected by the IceCube Neutrino Observatory

@@ Abstract
[…]
@@ 1 Introduction
It is widely believed that most cosmic-ray particles detected on Earth below EeV energy are produced within the Milky Way galaxy. Their energy spectrum and arrival directions are determined by the properties and distribution of their galactic sources and by their propagation through the turbulent and inhomogeneous interstellar magnetized plasma. Although cosmic-ray properties have been investigated for over a century, it was only in the last thirty years that extensive ground-based experiments have accumulated high-quality observations of the arrival direction distribution of cosmic rays (Nagashima et al., 1998; Hall et al., 1999; Guillian et al., 2007; Aglietta et al., 2009; Abdo et al., 2009; Munakata et al., 2010; De Jong, 2011; Aartsen, M. G. et al., 2016a; Amenomori et al., 2017; Bartoli et al., 2018; Abeysekara et al., 2018, 2019; Apel et al., 2019; Ahlers, 2019; He, 2023).
@@
[…]
@@ 2.1 Detector
The IceCube Neutrino Observatory (Aartsen, M. G. et al., 2017) is a cubic-kilometer-scale detector located at the geographic South Pole. The largest component of the detector is an in-ice array buried deep in the Antarctic ice shelf, consisting of 5,160 digital optical modules (DOMs), extending from a depth of 1,450 meters to about 2,450 meters. The in-ice array is augmented with a denser subarray called DeepCore (Abbasi et al., 2012b). IceCube also includes a surface array called IceTop (Abbasi et al., 2013), composed of tanks with DOMs inserted in transparent ice and designed to detect cosmic-ray showers above 100 TeV. This paper describes an investigation performed with the high-energy cosmic-ray muon data collected by the IceCube in-ice array.
@@
The IceCube in-ice array is designed to detect high-energy neutrinos of astrophysical origin. Neutrinos are observed through the secondary charged particles generated by their interaction in the ice inside and outside the instrumented volume or in the bedrock below. Such secondary particles, being relativistic, induce Cherenkov photons that propagate through the ice and are detected by the DOMs. The PMT produces a current pulse whenever a Cherenkov photon ejects an electron from its cathode; these pulses are digitized and time-stamped by the DOM onboard electronics before being sent to a data acquisition system for trigger determination and subsequent event reconstruction. To identify a physical event and filter out background signals caused by the 500-Hz PMTs’ dark noise, neighboring DOMs on the same string must satisfy a local coincidence in time within ±$\pm 1\,\mu$s. A trigger is issued when eight or more DOMs satisfying the local coincidence criterion record photons within a ±$\pm 5\,\mu$s time window (Abbasi et al., 2009; Aartsen, M. G. et al., 2017).
All events that trigger the IceCube detector are reconstructed using a likelihood-based method that accounts for light propagation in the ice. This Single-Photo-Electron (SPE) reconstruction maximizes the likelihood constructed from the probability density function for the arrival times of single photons at the locations of activated optical modules to reconstruct the muon track (Ahrens et al., 2004). This SPE likelihood algorithm is seeded with the results of a fast linefit reconstruction that minimizes the sum of the squares of the distances between the track and the recorded pulses at each optical module (Aartsen, M. et al., 2014).
@@ 2.2 Experimental Data
This analysis uses twelve full years of data collected by the completed IceCube detector (referred to as IC86) from May 13, 2011, through May 12, 2023. The data correspond to the good run list established by IceCube’s online monitoring system (Aartsen, M. G. et al., 2017).
The rate of events collected by the IceCube in-ice array used in this work is between 2.0 and 2.3 kHz, following the yearly modulation of the stratospheric temperature over the Antarctic continent (Gaisser & Verpoest, 2021; Verpoest et al., 2024). The volume of data collected amounts to about 1 TB per day. The limited bandwidth of the communication satellites that provide connectivity to the South Pole sets a practical limit on the amount of information that can be stored and transferred from each event that activates the detector’s trigger. For this reason, we utilize a compact data storage and transfer (DST) format to store the reconstruction results of all recorded events performed in situ before transferring the data to the Northern Hemisphere.
@@ 2.3 Simulation Data
We propagate the muons generated by CORSIKA through the ice and simulate the in-ice array response. The detector simulation — including detailed photon propagation, PMT response, and acquisition electronics — uses IceCube’s IceTray framework (DeYoung, 2004) with an interface to read CORSIKA formatted files. This simulation uses only high-energy muons reaching the in-ice array; all other shower components not contributing to the physical event are omitted to reduce memory requirements and computation time. Finally, the triggered muon events are reconstructed and processed with the same pipeline used for experimental data. .
@@ 3.1 Constructing Arrival Direction Sky Maps
Numerous methods are used to measure the anisotropies in cosmic-ray arrival direction, depending on the experiment being used, including: Rayleigh amplitude method (Linsley, 1975; Antoni et al., 2004), time-scrambling (Alexandreas et al., 1993), direct integration (Atkins et al., 2003), a $\chi^{2}$ iterative method optimized for gamma-ray searches (Amenomori et al., 2005), East-West (Bonino et al., 2011), and the iterative maximum likelihood method developed by Ahlers et al. (2016). This last method was utilized in the combined analysis with HAWC in Abeysekara et al. (2019). It is designed to correctly recover the amplitude of large-scale features in the anisotropy that are attenuated as a result of the difference between the instantaneous and integrated field of view (FoV). For detectors located away from the poles, the instantaneous exposure varies during a 24 hr period and does not match the full daily exposure. However, as noted by Ahlers et al. (2016), since IceCube’s instantaneous FoV is identical to the time-integrated one, the iterative method is equivalent to direct integration or time scrambling. For this reason, in this analysis, we employed the time-scrambling method, as in previous IceCube analyses (Abbasi et al., 2012a; Aartsen, M. G. et al., 2016a), to provide a “reference sky map” representing the experiment’s response to an isotropic cosmic-ray flux. This reference sky map also accounts for possible detection biases (e.g., nonuniform exposure and gaps in the experiment’s livetime).
@@
where $N_{i}$ and $\langle N\rangle_{i}$ represent the number of events for pixel $i$ in the data and reference sky maps, respectively. The statistical significance of each pixel is estimated using the method developed by Li & Ma (1983). We smooth the data and the reference sky maps for visualization purposes to observe large- and small-scale anisotropies at high significance (see Sec. 3.3). We use a process known as top-hat smoothing, which assigns a value in each pixel equal to the sum of all pixels’ content within a smoothing radius. Our default smoothing radius is $20^{\circ}$, consistent with previous IceCube publications (Abbasi et al., 2011, 2012a; Aartsen, M. G. et al., 2016a). Analyses of the cosmic-ray arrival direction distributions are conducted on unsmoothed data and reference maps.
The Monte Carlo simulation data described in Sec. 2.3 indicate that the median angular resolution of the entire data sample is about $3^{\circ}$ (reaching $1^{\circ}$ above 100 TeV). However, it surpasses $20^{\circ}$ above a zenith angle of $50^{\circ}$ at 10 TeV, with a significant improvement at higher energy. The poor angular resolution at large zenith angles results from the increasing energy threshold of cosmic-ray primaries, as muons must travel increasingly longer distances through the ice. As a result, the number of hit optical sensors decreases at larger zenith angles for a given primary energy (Abbasi et al., 2012a; Abeysekara et al., 2019). For visualization purposes, sky maps are masked to omit events with reconstructed zenith angles above $65^{\circ}$, above which the event rate is negligible. On the other hand, smaller sky coverage is used for quantitative data analyses (see Sec. 3.4). This study does not attempt to select events with better angular resolution, as we aim to obtain the largest data volume to determine anisotropy structures down to approximately $6^{\circ}$ (see Sec. 3.5).
@@ 3.2 Constructing Energy Datasets
To study the energy dependence of the cosmic-ray anisotropy, we split the experimental data into several sets of increasing cosmic-ray energy. This procedure uses the simulation data described in Sec. 2.3. In earlier studies on anisotropy, we relied on the data gathered during the construction of IceCube, which limited our capacity to generate sufficient simulation data for each experiment’s configuration (see Aartsen, M. G. et al. (2016a) and references therein). However, we now benefit from exclusively using the final detector configuration for the entire time period used in this work.
@@
The primary cosmic ray energy estimation is done on a statistical basis using the number of DOMs triggered by each event ($N_{\mathrm{channel}}$) and the corresponding muon trajectory’s reconstructed zenith angle ($\theta_{\mathrm{reco}}$). Higher deposited energy in the experiment triggers more DOMs. This happens with high-energy muons and high-multiplicity muon bundles, which are associated with more energetic or heavier cosmic-ray primary particles. In addition, events with a larger zenith angle must penetrate more ice before passing through IceCube’s in-ice array. Inclined muons therefore have higher mean energy at the surface, corresponding to higher primary cosmic-ray energy (Aartsen, M. G. et al., 2016b). We use the simulation data described in Sec. 2.3 to determine the dependency between the experiment’s response and the cosmic-ray primary particle energy as a function of the reconstructed zenith angle.
@@
Using the procedure described in Aartsen, M. G. et al. (2016a), Monte Carlo simulated events were placed into a three-dimensional histogram with axes corresponding to the observed $\log_{10}(N_{\mathrm{channel}})$, $\cos(\theta_{\mathrm{reco}})$, and the logarithm of the true energy of the primary cosmic-ray particle (bin sizes of 0.044, 0.07, and 0.05, respectively). After the histogram is populated, we select the median true energy value for each bin, effectively transforming the histogram into a lookup table that displays median energy as a function of $\log_{10}(N_{\mathrm{channel}})$ and $\cos(\theta_{\mathrm{reco}})$. The table is smoothed with a B-spline function (Whitehorn et al., 2013) to avoid statistical artifacts resulting from the limited simulation at high energies. The end result is used to define nine ranges of median primary cosmic-ray energy, as shown in Fig. 1a. The bands in the figure provide a proxy for estimating the primary energy of events, based on the measured variables $N_{\mathrm{channel}}$ and $\theta_{\mathrm{reco}}$. The described energy selection method yields an energy distribution in each event set defined by a colored band with a median value that is fairly constant as a function of $\theta_{reco}$. To ensure that the median energy on the lowest energy bin is also constant as a function of zenith angle, we exclude the events populating the white area in the bottom corner of Fig. 1a. Since the anisotropy changes with energy (see Sec. 3.3), ensuring a uniform energy response across the FoV provides an unbiased measurement of the arrival direction distribution.
@@
Table 2 displays the number of events selected for each energy set used in the analysis. This corresponds to about 82% of the total number of events collected by the IceCube in-ice array (see Table 1), a reduction stemming from the limited FoV (see Sec. 3.1), the exclusion of the lower right corner ($\log_{10}(N_{\mathrm{channel}})$, $\cos(\theta_{\mathrm{reco}})$) region in Fig. 1a, and the selection $N_{\mathrm{channel}}>10$, which removes events that are harder to reconstruct. For all energy bins, we apply a reconstructed zenith angle cut of 65°65$\degree$65 °. The only exception is the lowest energy bin, at $E_{\mathrm{median}}=13$ TeV, where we apply a reconstructed zenith angle cut of 60°60$\degree$60 °, since for larger zenith angles, events fall below the $N_{channel}>10$ threshold. For the highest energy bin, we apply only a lower bound in the energy selection. The primary cosmic-ray energy distributions corresponding to each of the nine selected regions are shown in Fig. 1b, along with the associated median energy. Note that the energy distributions of these sets overlap significantly. However, the energy sets are statistically independent as each event is placed in only one of them. The overlap between the different distributions results mainly from two reasons. First, as in the case of the angular resolution (see Sec. 3.1), there is no attempt to optimize the muon energy resolution with dedicated event selections. The goal is to have the largest number of events possible while retaining a reasonable quality. Secondly, the energy deposited by the muon bundles has wide event-by-event fluctuations, and even with a more accurate muon energy estimation, its relation to the parent cosmic-ray particles would still be affected by large fluctuations.
By comparing the H3a (Gaisser, 2012), Polygonato (Jörg R. Hörandel, 2003), and GST (Gaisser et al., 2013) composition models, we determined that the uncertainty in the median energy from different cosmic-ray flux models is less than 444 % in $\log_{10}(E/\mathrm{GeV})$. The predicted atmospheric muon fluxes depend on the assumed primary cosmic-ray spectrum and composition and the hadronic models that describe these interactions. Discrepancies among these models lead to an overall value of 20–40% in the predicted muon flux within the energy range from 1 TeV to 1 PeV (Riehn et al., 2020).